# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

### v3

@@ 4.3 Harnessing Prior Knowledge from Pretrained Models via Optimal Transport
Considering a leaf group, there may be data classes with varying levels of overlap in the latent space. For example, Figure 5 provides statistical results regarding the relative position of classes in leaf groups ”plants” and ”four-legged mammals”. We can see that in the group of animals, the L2-Wasserstein distance between class representations of ”mouse” and ”otter” (39.30) is more significant than the one between ”mouse” and ”hamster” (26.05), meaning that the class representation of the second pair can overlap more than the first one. Although focusing on classes within the same leaf group helps improve the ability to recognize difficult-to-identify classes, we still treat all classes in that group equally. Thus, the algorithm may inadvertently ignore the pairs of classes that are easily confused and need to be further distinguished.
Moreover, pretrained models are extensively trained on large datasets, which results in their substantial generalization abilities. That is the reason why they are considered as good starting points for the adaptation to downstream tasks [8, 4, 17, 15, 35]. In this work, we propose to exploit the pretrained models from a novel perspective, which can comprehend the use of the label-based tree-like taxonomy during training. In particular, we take advantage of these models to obtain prior assumptions about the relationships between the class representations on the latent spaces, thereby elegantly intervening in better separating each pair of classes within a leaf group.
@@
Firstly, to extract the relationship between the classes, we compute the L2-Wasserstein distance of feature distributions of each class pair. Specifically, we denote $D_{c_{i}}^{\Phi}$ be the distribution of class $c_{i}$ on the latent space of the pretrained model $f_{\Phi}$ to distinguish it from the class distribution on prompted latent space $f_{\Phi,\bm{P}}$, as described in Section 4.2. The distribution $D_{c_{i}}^{\Phi}$ is obtained in the form of a Gaussian Mixture model at the end of the respective task, only once and then kept fixed. In this way, when $t$ tasks have arrived, we have the corresponding sets of distributions $\{D_{c}^{\Phi}\}_{c\in Y_{1,t}}$ of all $m_{t}$ classes so far. Therefore, we gradually complete the WD-based matrix between pairs of classes:
@@
The fact is that the smaller the $W_{2}$ between the representations of two classes, the closer they are, and the more force can be needed to separate them. Therefore, we employ these results to obtain the corresponding weight factors, which control how much constraint needs to be applied to each pair. In particular, we compute the weight matrix:
$\Gamma=\left[\gamma_{ij}\right]_{m_{t}\times m_{t}}=\left[1/{\exp(M_{ij}/\delta)}\right]_{m_{t}\times m_{t}}$ (3)
where $\delta$ is a temperature. We then apply this information to obtain a weighted version of $\mathcal{L}_{g}$, in which the closer the two class distributions are, the larger the weight assigned, and they will be focused to push away. Consequently, our regularization loss becomes:
$\mathcal{L}_{g}(\cdot)=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}\gamma_{y_{x}y_{\tilde{x}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$ (4)
This strategy is completely economical and aligns well with the CL learning scheme, as the distance between representations of each class pair only needs to be computed once, and the matrix $M$ is continuously expanded when new classes arrive. Practically, when learning a new task, the first epoch is for capturing information about the behavior of the pretrained model on the data of this task. Moreover, this approach is also aligned with the findings in Cognitive Science [34, 2], showing that the accumulated experiences from past learning create momentum for learning new skills more effectively.
Finally, our objective function during training can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{CE}+\alpha\mathcal{L}_{g}+\beta\mathcal{L}_{all}$ (5)
@@ 5.1 Experimental setup
Baselines. We compare ours with the latest state-of-the-art prompt-based methods, including the methods using shared prompts for all tasks: L2P [49], DualPrompt [48], OVOR [18]; and the methods dedicates a distinct prompt set for each task like: S-Prompt++ [47], CODA-Prompt [38], HiDe-Prompt [44], CPP [25] (see Appendix A.2).
Metrics. We use two main metrics, including the Final Average Accuracy (FAA) and the Final Forgetting Measure (FFM), denoting the average accuracy and forgetting of all tasks after learning the sequence of tasks, respectively (see Appendix A.2). Besides, the detailed implementation is described in Appendix A.4.
(a) The role of $\mathcal{L_{G}}$
(b) The role of OT-based strategy
(c) Varying the value of $\alpha,\beta$
Figure 6: Ablation studies about our training strategy.
@@ 5.2 Experimental result
Table 1 presents the overall performance comparison between our proposed method and all the baselines. The key observation is that our method is the strongest one, by the gap between our method and the runner-up method being around 1.5% in terms of FAA on all considered datasets. Additionally, the results show that our method avoids forgetting better than all baselines, by the gap up to 17% on 5-Datasets.
Ablation studies about our training strategy
Figure 6 reports the ablation studies demonstrating the effectiveness of our training strategy. Particularly, compared to training tasks independently using Cross Entropy loss $\mathcal{L}_{CE}$ like in DualP, L2P, and CODA-P, exploiting the relationships between data classes with the label-based hierarchical taxonomy and the WD-based cost matrix helps improve FAA by about 5% to 10% (Figure 6(a)). Besides, when examining the role of exploiting additional prior information from pretrained backbones using the OT approach, we see that FAA is improved from 0.6% to 0.8% (Figure 6(b)). These results demonstrate the positive impact of this component, confirming the importance of exploiting correlations between class data during training. In both figures, the improvements on Split-CIFAR100 and 5-Datasets are the lowest, while it is more pronounced on Split-CUB-200. This may be because the groups of these two datasets (Split-CIFAR100 and 5-Datasets) have fewer overlapping classes, as the classes in each group likely have more recognizable features. Meanwhile, Split-CUB-200 is a dataset of birds with images that can be difficult for human eyes to recognize, thus so our method performs better.
In addition, Figure 6(c) provides the experimental results on the dataset Split-CUB-200, when varying $\alpha$ and $\beta$, which control the intensification of impact on each leaf group of $\mathcal{L_{G}}$ during training. The data shows that with a large enough value of $\beta$, our TCL is not sensitive to $\alpha$ within its acceptable range. Conversely, if $\alpha$ is small, the quality of the model can change more significantly.
t-SNE visualization of representation on latent space
Furthermore, Figure 4 illustrates the effect of our method in improving the model’s representation learning on Split-CIFAR100. Specifically, the classes are better clustered, and the separation between them is more distinct. Especially, the classes ’oak tree,’ ’willow tree,’ and ’pine tree’ are divided into clear clusters, rather than being mixed together as in the traditional training strategy, where tasks are trained independently.
@@
Table 2: Comparison when using different pretrained backbones
@@ 6 Conclusion and Discussion
In this work, we demonstrate the importance of organizing and exploiting data meaningfully rather than lumping it together for training. In particular, we propose organizing data into a tree-like taxonomy based on their label information, helping to identify groups of class data that are likely to confuse models. In this way, we encourage the models to focus and build deeper knowledge about each group, reducing forgetting and motivating more effective learning in subsequent tasks. Additionally, we propose leveraging the initial behavior of pretrained models to unveil a hidden structure of training data, providing an additional perspective to improve performance further. Finally, experimental results demonstrate the effectiveness of these components and our superiority over state-of-the-art baselines.
Limitations.
Despite this novel approach, the quality of the hierarchical taxonomy depends on the quality of expert knowledge. For example, if similar image classes are not assigned to the same leaf group in this label-based taxonomy, the constraint we put on each such group may not perform as expected.
Supplement to “Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning“
@@ A.4 Implementation Details
Appendix B Using ChatGPT to build tree-like taxonomy during a sequence of tasks incrementally
@@
"Four-legged": ["leopard", "rabbit", "mouse", "camel"]
"Aquatic": ["trout", "aquarium_fish"],
@@
"Four-legged": ["leopard", "rabbit", "mouse", "camel"]
"Aquatic": ["trout", "aquarium_fish"],
@@
"Four-legged": ["leopard", "rabbit",
"mouse", "camel", "otter"]
"Aquatic": ["trout", "aquarium_fish",
"shark", "seal", "lobster"],
@@
"Buildings": ["skyscraper", "house"]
T3 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": ["leopard", "rabbit", "mouse",
"camel", "otter", "chimpanzee", "squirrel"]
"Aquatic": ["trout", "aquarium_fish", "shark",
"seal", "lobster", "dolphin", "flatfish", "crab"],
"Reptiles": ["snake"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower"],
"Containers": ["bottle", "bowl"],
"Appliances": ["television"]
"Vehicles": {
"Wheeled": ["motorcycle"]
"Structures": {
"Buildings": ["skyscraper", "house"]
T4 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew"
"Two-legged": ["woman"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl"],
"Appliances": ["television"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"]
T5 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox"
"Two-legged": ["woman"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle"]
"Plants": {
"Flowers": ["rose", "tulip"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"]
T6 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle"],
"Others": ["worm"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T7 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle", "lion"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal",
"lobster", "dolphin", "flatfish", "crab", "ray"
"Reptiles": ["snake", "lizard"],
"Insects": ["caterpillar", "beetle", "bee", "cockroach", "spider"],
"Others": ["worm"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest", "cloud", "plain"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T8 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle", "lion"
"Two-legged": ["woman", "man", "baby"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard", "turtle"],
"Insects": ["caterpillar", "beetle", "bee", "cockroach", "spider"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair", "couch", "table"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle", "pickup_truck", "tractor"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T9 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle",
"lion", "tiger", "wolf", "beaver", "possum", "skunk"
"Two-legged": ["woman", "man", "baby", "boy"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard",
"turtle", "crocodile", "dinosaur"],
"Insects": ["caterpillar", "beetle",
"bee", "cockroach", "spider"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower", "orchid"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock"],
"Furniture": ["bed", "chair", "couch", "table"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle",
"pickup_truck", "tractor"],
"Air": ["rocket"],
"Rail": ["train"]
"Structures": {
"Buildings": ["skyscraper", "house"],
"Bridges": ["bridge"],
"Others": ["road"]
T10 = {
"Natural": {
"Animals": {
"Mammals": {
"Four-legged": [
"leopard", "rabbit", "mouse", "camel", "otter",
"chimpanzee", "squirrel", "porcupine", "shrew",
"hamster", "raccoon", "fox", "kangaroo", "cattle",
"lion", "tiger", "wolf", "beaver", "possum", "skunk",
"elephant", "bear"
"Two-legged": ["woman", "man",
"baby", "boy", "girl"]
"Aquatic": [
"trout", "aquarium_fish", "shark", "seal", "lobster",
"dolphin", "flatfish", "crab", "ray", "whale"
"Reptiles": ["snake", "lizard",
"turtle", "crocodile", "dinosaur"],
"Insects": ["caterpillar", "beetle",
"bee", "cockroach", "spider", "butterfly"],
"Others": ["worm", "snail"]
"Plants": {
"Flowers": ["rose", "tulip", "poppy", "sunflower", "orchid"],
"Trees": ["palm_tree", "willow_tree",
"maple_tree", "pine_tree", "oak_tree"],
"Fruits": ["sweet_pepper", "apple", "pear", "orange"],
"Fungi": ["mushroom"]
"Environment": {
"Natural Features": ["mountain", "forest",
"cloud", "plain", "sea"]
"Man-Made": {
"Objects": {
"Tools": ["lawn_mower", "tank"],
"Containers": ["bottle", "bowl", "plate", "can", "cup"],
"Appliances": ["television"],
"Instruments": ["keyboard", "lamp", "clock", "telephone"],
"Furniture": ["bed", "chair", "couch", "table", "wardrobe"]
"Vehicles": {
"Wheeled": ["motorcycle", "bicycle",
"pickup_truck", "tractor", "bus"],
"Rail": ["train", "streetcar"],
"Air": ["rocket"]
"Structures": {
"Buildings": ["skyscraper", "house", "castle"],
"Bridges": ["bridge"],
"Others": ["road"]

### v4

@@ 4.3 Harnessing Prior Knowledge from Pretrained Models via Optimal Transport
Considering a leaf group, there may be data classes with varying levels of overlap in the latent space. For example, Figure 5 provides statistical results regarding the relative position of classes in leaf groups ”plants” and ”four-legged mammals” in the latent space of a pretrained model. We can see that in the group of animals, the L2-Wasserstein distance between class representations of ”mouse” and ”otter” (39.30) is more significant than the one between ”mouse” and ”hamster” (26.05), meaning that the class representation of the second pair can overlap more than the first one. This empirical result suggests that although the strategy in Sec 4.2 helps improve the model’s ability to recognize difficult-to-identify classes, we still treat all classes in that group equally. Thus, the algorithm may inadvertently ignore the pairs of classes that are easily confused and need to be further distinguished.
Figure 5: L2-Wassertein distance between classes (Split-CIFAR-100) in latent space of a pretrained backbone (Sup-21K). Within a leaf group (i.e., ”plant” or ”four-legged mammals”), there may be data classes with varying levels of correlation in the latent space.
Moreover, pretrained models are extensively trained on large datasets, which results in their substantial generalization abilities. That is the reason why they are considered as good starting points for the adaptation to downstream tasks [10, 4, 18, 38]. In this work, we propose utilizing the pretrained models from another novel perspective, which can comprehend the use of the label-based taxonomy (Sec. 4.2) during training. Particularly, we take advantage of these models to obtain prior assumptions about the relationships between the class representations, thereby elegantly intervening in better separating each pair of classes within each leaf group.
@@
More specifically, to extract the relationship between the classes, we compute the L2-Wasserstein distance of feature distributions of each class pair. We denote $D_{c_{i}}^{\Phi}$ be the distribution of class $c_{i}$ on the latent space of the pretrained model $f_{\Phi}$ to distinguish it from the one on prompted latent space $f_{\Phi,\bm{P}}$, as described in Sec. 4.2, Line 364. The distribution $D_{c_{i}}^{\Phi}$ is obtained in the form of a Gaussian Mixture model at the end of the respective task, only once and then kept fixed for the next reuse; therefore, it does not impose a computational burden. In this way, when $t$ tasks have arrived, we have the corresponding sets of distributions $\{D_{c}^{\Phi}\}_{c\in Y_{1,t}}$ of all $m_{t}$ classes so far. Thus, we can gradually complete the WD-based matrix between pairs of classes:
@@
The fact is that the smaller the $W_{2}$ between the representations of two classes, the closer they are, and the more force can be needed to separate them. Thus, we employ these results to obtain the corresponding weight factors, which control how much constraint needs to be applied to each class pair. Particularly, we compute the weight matrix:
$\Gamma=\left[{\color[rgb]{0,0,1}\bm{\gamma_{ij}}}\right]_{m_{t}\times m_{t}}=\left[1/{\exp(M_{ij}/\delta)}\right]_{m_{t}\times m_{t}}$ (4)
where $\delta$ is a temperature. We then apply this information to obtain a weighted version of $\mathcal{L}_{g}$, in which the closer the two class distributions are, the larger the weight assigned, and they will be further focused to push away:
$\mathcal{L}^{{}^{\prime}}_{g}(\cdot)=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{1..t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{1..t}}{\color[rgb]{0,0,1}\bm{\gamma_{y_{x}y_{\tilde{x}}}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$ (5)
This strategy is completely economical and aligns well with the CL learning scheme, as the distance between representations of each class pair only needs to be computed once, and the matrix $M$ is continuously expanded when new classes arrive. Practically, when learning a new task, the first epoch is for capturing information about the behavior of the pretrained model on the data of this task. Moreover, this approach is also aligned with the findings in Cognitive Science [37, 2], showing that the accumulated experiences from past learning create momentum for future learning.
Finally, the final objective function of our full method can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{CE}+\alpha\mathcal{L}^{{}^{\prime}}_{g}+\beta\mathcal{L}_{all}$ (6)
@@ 5.1 Experimental setup
Baselines. We compare ours with the various state-of-the-art prompt-based methods, including L2P [52], DualPrompt [51], OVOR [20], S-Prompt++ [50], CODA-Prompt [41], HiDe-Prompt [47], CPP [27] (see Appendix A.2).
Metrics. We use two main metrics, including the Final Average Accuracy (FAA) and the Final Forgetting Measure (FFM) (see Appendix A.2). Besides, the detailed implementation is described in Appendix A.4.
(a) The role of our $\mathcal{L_{G}}$ Eq (1)
(b) The role of our OT technique (Sec 4.3) in $\mathcal{L_{G}}$
(c) Varying the value of $\alpha,\beta$ (Split-CUB-200)
Figure 6: Ablation studies about our training strategy. Figure (b) depicts the performance improvement after applying our OT-based technique to our taxonomy strategy. The efficiency of this technique w/o our taxonomy strategy is presented in Appendix C.3.
@@ 5.2 Experimental result
Table 1 presents the overall performance comparison between our proposed method and other baselines. The key observation is that our method is the strongest one, because the gap compared with the runner-up method being around 1.5% of FAA on all the datasets. Additionally, our method avoids forgetting better than all baselines, by the gap up to 17% on 5-Datasets.
Ablation studies.
Figure 6 reports the ablation studies of our training strategy. Particularly, compared to training tasks independently using Cross Entropy loss $\mathcal{L}_{CE}$ like in DualP and L2P, exploiting the relationships between data with the label-based taxonomy and the OT-based strategy (ours) helps improve FAA by about 5% to 10% (Figure 6(a)). Besides, when examining the role of exploiting additional prior information from pretrained backbones using the OT approach upon the taxonomy, we see that FAA is improved from 0.6% to 0.8% (Figure 6(b)). These results demonstrate the positive impact of this component. In both figures, the improvements on Split-CIFAR100 and 5-Datasets are the lowest, while it is more pronounced on Split-CUB-200. This may be because the groups of these two datasets (Split-CIFAR100 and 5-Datasets) have fewer overlapping classes, as the classes in each group likely have more recognizable features. Meanwhile, Split-CUB-200 is a dataset of birds with images that can be difficult for human eyes to recognize, thus so our method performs better. Furthermore, Figure 6(c) provides the experimental results when varying $\alpha$ and $\beta$. The data show significant degradations of model performance when each loss function of $\mathcal{L_{G}}$ is eliminated, demonstrating their specific roles. In particular, w/o using our taxonomy and OT-based strategy ($\alpha\mathcal{L}_{g}=0$), the performance can decrease up to 5.65%. Besides, the effect of $\mathcal{L}_{all}$ can be clearly seen when $\beta=0$, as it helps constrain the relative correlations of all classes.
t-SNE visualization on latent space.
Figure 4 illustrates the effect of our method in improving the model’s representation learning on Split-CIFAR100. Specifically, the classes are better clustered, and the separation between them is more distinct. Especially, the classes ”oak tree”, ”willow tree”, and ”pine tree” are divided into clear clusters, rather than being mixed together as in the traditional training strategy, where tasks are trained independently.
The impact of $\delta$ in OT technique.
Figure 7 illustrates the dependence of model performance on the value of the coefficient $\delta$ in our OT technique (see Eq.4). The results show that if $\delta$ is too small, the performance can go down significantly, even worse than HiDE-Prompt. This occurs because the values of the weight matrix will be too large, leading to classes, which have strong correlations in the latent space of $f_{\Phi}$, being pushed too far apart. This causes an imbalance compared to the overall correlation of all classes and unexpected overlapping. Conversely, if $\delta$ is set too large, the corresponding weight will be small, preventing achieving optimal efficiency. For Split-CIFAR-100 and Split-ImageNet-R, the optimal values are $\delta=500$ and $100$, respectively.
Figure 7: Model performance when $\delta$, in Eq (4), varies.
@@
Table 2: Comparison when using different pretrained backbones.
Other results.
We also provide experimental results in Appendix sections, related to model performance when using different taxonomy structures, and analyze other aspects of our OT technique, the coefficients $\alpha,\beta,\tau$ when using loss functions. Additionally, we verify the correlation between taxonomy and visual space, as well as discuss the current limitations and potential impacts of the proposed method. See Appendix C.
@@ 6 Conclusion and Discussion
In this work, we demonstrate the importance of meaningfully organizing data information rather than just lumping them together for training. Particularly, we propose arranging data labels into tree-like taxonomies to identify data groups that are likely to confuse models. This approach encourages the models to focus and develop deeper knowledge about each group, reducing forgetting and motivating more effective learning in subsequent tasks. Additionally, we propose leveraging the initial behavior of pretrained models to obtain hidden structures of training data, providing a new perspective to further enhance performance. Finally, the experimental results demonstrate our superiority over state-of-the-art baselines.
Supplementary Material
@@ A.4 Implementation Details
Appendix B Using LLMs (i.e., ChatGPT, DeepSeek, etc.,) to build tree-like taxonomy during a sequence of tasks, incrementally
@@
"Four-legged": ["leopard",
"rabbit", "mouse", "camel"]
"Aquatic": ["trout",
"aquarium_fish"],
@@
LLMs used to build taxonomies Split CIFAR-100 Split Imagenet-R Split CUB-200
Sup-21K iBOT-21K Sup-21K iBOT-21K Sup-21K iBOT-21K
Baseline (HiDE-Prompt) 92.61 93.02 75.06 70.83 86.56 78.23
Llama-3-70b-Groq 93.58 93.82 77.01 73.12 88.00 79.05
GPT-4o-Mini 93.94 94.01 76.41 72.12 87.93 79.02
Gemini-1.5-Pro 93.35 93.43 75.76 71.42 87.02 78.43
DeepSeek-R1 93.56 93.62 76.24 71.93 87.65 78.72
Mistral-Medium 93.12 93.35 75.54 71.12 87.00 78.22
Claude-3.5-Haiku 93.65 93.62 76.12 71.98 87.83 78.75
Average results of LLMs $93.53_{\pm 0.28}$ $93.64_{\pm 0.24}$ $76.18_{\pm 0.52}$ $71.95_{\pm 0.69}$ $87.57_{\pm 0.45}$ $78.8_{\pm 0.33}$
Table 3: Performance comparison when using different LLMs to generate the corresponding taxonomy
@@
"Four-legged": ["leopard",
"rabbit", "mouse", "camel"]
"Aquatic": ["trout",
"aquarium_fish"],
@@
"Four-legged": ["leopard",
"rabbit", "mouse", "camel",
"otter"]
"Aquatic": ["trout",
"aquarium_fish",
"shark", "seal",
"lobster"],
@@
"Buildings": ["skyscraper",
"house"]
T3 = ...